OpenStudy (anonymous):

what is 3i(1-i) in a+bi form??? I got 3i-3

OpenStudy (whpalmer4):

What is \(3i*(-i)=\) Remember, \(i^2=-1\)

OpenStudy (anonymous):

-3i(squared) so it's 3???

OpenStudy (anonymous):

^yes

OpenStudy (whpalmer4):

Yes. So your new, improved answer is?

OpenStudy (anonymous):

3i??? Im sorry this is a hard topic for me am I no right???

OpenStudy (whpalmer4):

What is \(3i*1=\) What is \(3i*-i=\) Add them together for your answer

OpenStudy (whpalmer4):

It's just like calculating \(3x(1-x)\) except that after you do so, you replace \(x^2\) with \(-1\)

OpenStudy (anonymous):

I mean I got 0i

OpenStudy (anonymous):

is that wrong???

OpenStudy (whpalmer4):

Yes, that is incorrect. \[3i(1-I)=3i*1-3i*i=3i-3i^2=3i-3(-1)=\]

OpenStudy (whpalmer4):

First bit should have been \[3i(1-i)\] but iOS "helpfully" capitalizes it for me...

OpenStudy (anonymous):

I got 3i-3 again

OpenStudy (whpalmer4):

Or if you prefer \[3i(1-i)=3(i*1-i*i)=3(i-i^2)=3(i-(-1))=3i+3\] If you are still getting something else, you need to review manipulation of negative numbers.

OpenStudy (anonymous):

I got 3i-3 again

OpenStudy (anonymous):

okay thanks i'll just go review with my teacher

OpenStudy (whpalmer4):

-1*-3=

OpenStudy (anonymous):

3

OpenStudy (whpalmer4):

Then why aren't you doing that in this problem?!?!?!

OpenStudy (anonymous):

I don't know but okay

OpenStudy (whpalmer4):

Look carefully at what I posted

OpenStudy (whpalmer4):

-(-1)=+1

OpenStudy (anonymous):

okay

OpenStudy (whpalmer4):

-(-3)=+3

OpenStudy (anonymous):

yes I understand that

OpenStudy (whpalmer4):

Then you should be getting the right answer if you pay attention to what you are doing. Can you write out all of your steps for me?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

One line t a time, so I can stop you right away

OpenStudy (anonymous):

so first I distributed the 3i to 1 and -i... I got 3i-3i^2

OpenStudy (whpalmer4):

Yes, go ahead

OpenStudy (anonymous):

then I substituted the i^2 to -1

OpenStudy (whpalmer4):

Show me your work.

OpenStudy (whpalmer4):

I don't know why it is that so many people on OpenStudy, when asked to show their work choose instead to describe it. If we are trying to spot your mistake, we want to see exactly what you did, right?

OpenStudy (anonymous):